OpenStudy (anonymous):

Help Please? How do I simplify the expression -5+i/2i

OpenStudy (math&ing001):

Multiply by "i" to get rid of it from the denominator.

OpenStudy (anonymous):

ok so now it is -5+i/2

OpenStudy (math&ing001):

Like this : \[\frac{ -5+i }{ 2i }*\frac{ i }{ i }=\frac{ -5i-1 }{ -2 }\] You can simplify it further.

OpenStudy (math&ing001):

I meant multiply the nominator and the denominator by "i".

OpenStudy (anonymous):

oh ok that makes a little more sense...I think I got it now. Thank you!

OpenStudy (math&ing001):

Welcome ^^